OpenStudy (anonymous):

A= a + b + c/ 3 To solve for b in this formula, the first step is to: A.) Subtract 3 from both sides of the equation. B.) Divide both sides of the equation by 3. C.) Multiply both sides of the equation by 3. D.) Add a -3 to both sides of the equation.

OpenStudy (anonymous):

I would say C

OpenStudy (anonymous):

opposite of addition is subtraction

OpenStudy (anonymous):

You will have to get rid of 3 that is in the denominator..

OpenStudy (anonymous):

So... A?

OpenStudy (anonymous):

Nope..

OpenStudy (anonymous):

A= a + b + c/ 3 Multiply 3 on both sides 3A=a+b+c Subtract a and c on both sides 3A-a-c=b

OpenStudy (anonymous):

On Right Hand Side, 3 is in division.. What is the opposite of Division??

OpenStudy (anonymous):

subtraction?

OpenStudy (anonymous):

@InsanelyChaotic you said earlier opposite of addition is subtraction So the opposite of division is multiplication.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Opposite of Division is Multiplication and vice-versa.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Yep.. Good..